A florist wishes to make bouquets of mixed spring flowers. Part of each bouquet is to be tulips at $5 a bunch and part is to be daffodils at $2 a bunch. How many bunches of each should she use to make 15 bouquets which she can sell for $3 a bouquet.
Answer by macston(5194)

T=Tulips; D=Daffodils
T+D=15
T=15-D
$5T+$2D=$45 Substitute for T
$5(15-D)+$2D=$45
$75-$5D+$2D=$45 Subtract $75 from each side
-$3D=-$30 Divide each side by -$3
D=10 ANSWER 1: She should use 10 bunches of daffodils.
T=15-D=15-10=5 ANSWER 2: She should use 5 bunches of tulips
CHECK
5 Tulips cost $25, 10 Daffodils cost $20 for a total of $45 for 15 bunches ($3 per bunch).